# Enhancing the NAO: Extending Capabilities of Legacy Robots for Long-Term Research

### v1

@@ Abstract
Many research groups face challenges when legacy (unsupported) robotic platforms lose manufacturer support and cannot accommodate modern sensing, speech, and interaction capabilities. We present the Enhanced NAO, a revitalized version of Aldebaran’s NAO robot that uses upgraded microphones, RGB-D and thermal cameras, and additional compute resources in a fully self-contained package. This system combines cloud and local models for perception and dialogue, while preserving the NAO’s expressive body and behaviors. In a pilot validation study, the Enhanced NAO delivered significantly higher conversational quality and stronger user preference compared to the NAO AI Edition, without increasing response latency. Key upgrades, such as beamforming microphones and low-latency audio processing, reduced artifacts like self-hearing and improved multi-party separation. Expanded visual and thermal sensing established a foundation for future interaction capabilities. Beyond the NAO, our framework provides a platform-agnostic strategy for extending the lifespan and research utility of legacy robots, ensuring they remain valuable tools for human-robot interaction.
@@ I Introduction
The Aldebaran NAO [1] has been a staple in human–robot interaction (HRI) research for its approachable design, ease of use, and multimodal sensors. With Aldebaran’s recent bankruptcy and uncertainty around ongoing support, many labs risk losing a valuable research and education platform [2]. This challenge is not unique to NAO: Pepper was discontinued in 2021 [3], Rethink Robotics shut down in 2018, leaving Baxter obsolete [4, 5], and Willow Garage closed in 2014, ending support for the TurtleBot (first and second generations) and PR2 [6, 7, 8]. To address this broader problem of platforms becoming unsupported, we present a platform-agnostic framework for sustaining and enhancing obsolete robots, using the NAO as a central example.
Updating obsolete robots is challenging due to closed designs and outdated onboard compute resources. For NAO, prior work has added external computers [9], updated the operating system, [10], or integrated cloud services [11], but these approaches trade off portability, scope, or robustness.
We address these challenges with a fully integrated upgrade package (Fig. 1), featuring dual head-mounted cameras for multimodal perception, a Raspberry Pi Compute Module 5 (CM5) [12], a Seeed Studio ReSpeaker 4 Mic Array v2.0 [13] for robust speech pipelines, and a dedicated battery for peripheral power. This modular design extends sensing, processing, and interaction capabilities without reliance on manufacturer support or tethering.
Our contributions are threefold: (1) a complete hardware design, (2) an integration framework, and (3) supporting resources to be released open-source after review to enable widespread adoption. In a pilot validation study, the Enhanced NAO (with the custom upgrade package) demonstrated richer verbal interaction (e.g., conversation quality, participant preference) without added latency compared to the NAO AI, while expanded sensing and processing functionalities establish a practical blueprint for sustaining and enhancing legacy robots.
Figure 1: Enhanced NAO robot with integrated sensing, computing, and power systems. A. Back isometric view showing the ReSpeaker, custom CM5 carrier board, Hailo-8L, and portable battery pack. B. Front view highlighting the Orbbec Gemini 2 L depth camera, Seek thermal camera (S314SPX), and custom peripheral breakout board.
@@ II Related Work
Researchers have extended the life of legacy research platforms (e.g., NAO [1], Baxter [4], PR2 [7], first and second generation TurtleBots [6]) through added compute resources [9, 14, 15], updated sensors [14, 16, 17, 18, 19, 20, 21], modified end-effectors [22, 17, 23], and alternative software stacks [10, 24, 25, 26, 27, 28, 29, 21].
For the NAO, researchers pursued software and hardware strategies[10, 24, 9, 14]. Software efforts replaced the proprietary NAOqi head-unit with Ubuntu [24] and ROS2 [10], but these require flashing new firmware[24] and are limited by the NAO’s small eMMC, requiring a permanently attached USB stick for operation [24]. Hardware approaches preserved the factory firmware but added external compute and peripherals via short, wired connections [9, 14] , enabling capabilities not supported by the stock head unit.
Similar modifications exist for other humanoid and non-humanoid robots. Prior software upgrades have replaced ROS1 with ROS2 for Baxter [25], abstracted low-level ROS2 interfaces for PR2 [27], and introduced new inverse kinematic libraries for smoother control of both robots [26, 28]. Hardware upgrades have included custom end-effectors [22, 17, 23, 30], clothing and tactile sensors to improve human-robot hugs [19], and RGB-D cameras for teleoperation and object detection [16, 17, 18]. Non-humanoid robotic platforms, such as the early TurtleBots have received comparable software [20, 29] and hardware upgrades [15, 21], extending their utility beyond their official support.
@@ III Materials and Methods
Appropriateness ($A$) Flat tone Moderately engaging tone Warm, polite, and engaging tone
@@ III-A NAO Robot
We use the NAO v6 humanoid robot [1], a 58 cm-tall, 5.6 kg bipedal platform with 25 degrees of freedom (DOF), powered by an Intel Atom E3845 CPU with 4 GB RAM and a 32 GB eMMC storage running a Gentoo‑based Linux OS. The robot has two head-mounted HD cameras, a torso-mounted inertial measurement unit (IMU), and two ultrasonic sonars, four directional microphones, two speakers, nine capacitive touch sensors (three in the head, three per hand), and four bumper switches (two per foot). The platform supports Ethernet, Wi-Fi, and USB. Its lithium‑ion battery provides approximately 60-90 minutes of operation.
The NAO is controlled over a socket interface exposed by the proprietary NAOqi framework. Aldebaran provided a C++ library, Python bindings, and the Choregraphe GUI, though these require a local network connection to communicate with the socket interface. While the NAO v6 supports local code execution, it is restricted to Python 2.7, limiting developer flexibility [10].
RobotLab further markets the NAO AI Edition (hereafter referred to as “NAO AI”) with added cloud-based conversational features[31]. In contrast, our work builds on the base NAO v6 with custom hardware and software upgrades, collectively referred to as the “Enhanced NAO:”
@@ III-B1 Computing and Power System
A Raspberry Pi CM5 serves as the main processor, featuring a Broadcom BCM2712 quad-core 64-bit Arm Cortex-A76 (Armv8) running at 2.4 GHz [12]. The CM5 is mounted on a carrier board with USB, a GPIO, and PCIe expansion, supporting a Hailo-8L accelerator for 13 TOPS of onboard vision-based inference [32].
Peripheral connections include USB ports and CSI/DSI interfaces for additional sensors (e.g., Seek Thermal Camera). The system is powered by a 5 V/5 A USB-C battery pack that sustains all peripherals. Mechanical enclosures, including backpack and sensor housings, were 3D printed, with mounts adapted from open-source designs [14] and custom components, as needed.ubsubsectionAudio System Audio capture is provided by the ReSpeaker with an embedded XMOS XVF-3000, offering built-in echo cancellation, noise suppression, and direction-of-arrival estimation [13].
III-B2 Camera System
The camera sensing suite integrates an Orbbec Gemini 2 L depth camera [33] and a Seek Thermal camera (S314SPX) [34]. The Orbbec provides synchronized RGB-D streams with native OpenCV compatibility, while the Seek Thermal delivers thermography data and RGB thermal visualizations.
Figure 2: Sample camera frames produced by the enhanced camera system. A) RGB and B) depth frames from the Orbbec Gemini 2 L and (C) thermal frame from the Seek Thermal (S314SPX) camera.
@@ III-C Software upgrades
The CM5 runs Raspberry Pi OS with system nodes implemented in Python and C++.
@@ III-C1 Camera Fusion
The CM5 processes RGB, depth, and thermal streams locally. Frames are synchronized and passed through the Hailo-8L for human detection, face region extraction, and landmark identification in thermal and RGB data. This pipeline establishes integration of thermal and visual cues for future interaction logic. Although not used in our pilot study, we implemented a demonstration pipeline integrating the Orbbec Gemini 2 L camera, Seek Thermal module, and the Hailo-8L (Fig. 2).
@@ III-C2 Verbal Interaction Pipeline
Figure 3: Audio processing pipeline, starting with audio input via the ReSpeaker and ending with response audio played back through the NAO speakers.
The audio pipeline (Fig. 3) begins with the ReSpeaker, which outputs raw, processed, and echo-canceled streams. These signals are processed by the Open Embedded Audition System (ODAS) [35], for real-time speaker separation and sound source localization.
Deepgram [36] handles speech recognition, bypassing NAO’s built-in speech-to-intent engine to enable unconstrained dialogue. After transcription, LLM agents perform three steps: a) detecting whether the user is speaking, b) cleaning transcription errors, and c) identifying turn completion to refine the input before a final LLM agent generates the reply.
To reduce latency, responses are streamed immediately using Amazon Polly [37]. Because Polly does not natively support streaming, we developed a sentence-segmentation method that sends ordered requests, enabling a “think and speak” behavior and reducing the delay between user input and robot output.
@@ III-D Pilot validation study
Figure 4: A participant engaging in a multi-turn conversation with the Enhanced NAO during the pilot validation study.
We conducted a within-subjects, counter-balanced pilot study with eight participants (four male presenting, four female presenting) to compare interactions with the NAO AI and the Enhanced NAO. Each 15-minute session included two conditions, with no compensation provided. In each condition, participants held a short conversation ($\geq 4$ turns; turn defined as a user utterance followed by a robot response). As shown in Fig. 4, the NAO was placed on a table facing the seated participant, bringing it closer to eye level. To ensure consistency, the first turn was always a user greeting to the robot; subsequent turns were open-ended, with the option to prompt the robot for a joke, if needed. After both conversations, participants indicated which interaction they preferred (e.g., in terms of ease and naturalness of conversation).
@@ III-E Quantifying quality of responses
We assessed conversational quality using a two-level framework adapted from LLM evaluation methods [38]. The original framework included coherence, accuracy, clarity, relevance, and efficiency; we adapted it for conversational contexts by replacing accuracy with specificity, removing efficiency, and adding appropriateness and timing. This framework allows us to capture objective response quality and subjective user experience.
@@
where $W_{r}$ and $W_{u}$ denote the number of words spoken by the robot and the user, respectively. Their ratio quantifies relative talkativeness. Using Equation 3, an $R=1$ indicates balanced turn lengths, $R>1$ indicates robot-dominated speech, and $R<1$ indicates user-dominated speech.
@@ IV Results
Feature NAO AI [31] Enhanced NAO[1] with Hardware/Software Upgrades
@@
Microphones Four omnidirectional microphones; four channels at 48kHz or one channel at 16kHz. Accessible via NAOqi ALAudioDevice in 170ms chunks [1]. No algorithms to support ASR pipelines. Replaces existing NAO microphones
Seeed Studio ReSpeaker 4 Mic Array [13]
@@
Speech (TTS) Converts via NAOqi ALTextToSpeech. Supports pitch, speed, volume, pause, emphasis control, and voice effects. [39] Flexible with NAO’s default TTS engine, local TTS engines, or cloud-based TTS services. Tested implementation uses Amazon Polly [37].
Speech (ASR) Via NAOqi ALSpeechRecognition. Only recognizes predefined keywords. Requires a set vocabulary and returns matches with confidence scores. Supports word spotting within longer speech. [40] Flexible transcription using a dedicated ASR channel from the ReSpeaker. Tested implementation uses Deepgram to facilitate real-time speech-to-text.
@@
Static, intent-driven system focused on matching predefined intents with spoken keywords that trigger pre-determined responses. TTS engine has limited voices and lacks dynamic range. Supports open-ended dialogue with Deepgram [36], uses prompt engineering to tune an LLM [41] to generate relevant responses, and supports dynamic TTS voice generation via Amazon Polly [37].

### v2

@@ Abstract
Legacy (unsupported) robotic platforms often lose research utility when manufacturer support ends, preventing integration of modern sensing, speech, and interaction capabilities. We present the Enhanced NAO, a revitalized version of Aldebaran’s NAO robot featuring upgraded beamforming microphones, RGB-D and thermal cameras, and additional compute resources in a fully self-contained package. This system combines cloud-based and local models for perception and dialogue, while preserving the NAO’s expressive body and behaviors. In a pilot user study validating conversational performance, the Enhanced NAO delivered significantly higher conversational quality and elicited stronger user preference compared to the NAO AI Edition, without increasing response latency. The added visual and thermal sensing modalities established a foundation for future perception-driven interaction. Beyond this implementation, our framework provides a platform-agnostic strategy for extending the lifespan and research utility of legacy robots, ensuring they remain valuable tools for human-robot interaction.
@@ I Introduction
The Aldebaran NAO [aldebaran_nao_2025] has been a staple in human–robot interaction (HRI) research for its approachable design, ease of use, and multimodal sensors. With Aldebaran’s recent bankruptcy and uncertainty around ongoing support, many labs risk losing a valuable research and education platform [strathearnSobolewska2025nao]. This challenge is not unique to NAO: Pepper was discontinued in 2021 [nussey2021softbank], Rethink Robotics shut down in 2018, leaving Baxter obsolete [lawrenceCarol2019rethink, rethink2012baxter], and Willow Garage closed in 2014, ending support for the TurtleBot (first and second generations) and PR2 [garage2011turtlebots, garage2012pr2, willow_garage]. To address this broader problem of platforms becoming unsupported, we present a platform-agnostic framework for sustaining and enhancing obsolete robots, using the NAO as a central example.
Updating obsolete robots is challenging due to closed designs and outdated onboard compute resources. For NAO, prior work has added external computers [mattamalaNaoBackpack2017], updated the operating system, [bono2024openaccessnaooan], or integrated cloud services [gestrin2024NaoChat], but these approaches trade off portability, scope, or robustness.
We address these challenges with an integrated upgrade package (Fig. 1), featuring dual head-mounted cameras for multimodal perception, a Raspberry Pi 5 [rpi5_2025], a Seeed Studio ReSpeaker 4 Mic Array [respeaker_2023] for robust speech pipelines, and a dedicated battery for peripheral power. This modular design extends sensing, processing, and interaction capabilities without reliance on manufacturer support or tethering.
Here, we contribute: (1) a complete hardware design, (2) an integration framework, and (3) supporting resources available upon request to enable widespread adoption. In a pilot validation study, the Enhanced NAO (with the custom upgrade package) demonstrated richer verbal interaction (e.g., conversation quality, participant preference) without added latency compared to the NAO AI, while expanded sensing and processing functionalities establish a practical blueprint for sustaining and enhancing legacy robots.
Figure 1: Enhanced NAO robot with integrated sensing, computing, and power systems. A. Back isometric view showing the ReSpeaker 4 Mic Array, Raspberry Pi 5, Hailo-8L, and battery pack. B. Front view highlighting the Orbbec Gemini 2 L depth camera and Seek Thermal camera (S314SPX).
@@ II Related Work
Researchers have extended the life of legacy research platforms (e.g., NAO [aldebaran_nao_2025], Baxter [lawrenceCarol2019rethink], PR2 [garage2012pr2], TurtleBots [garage2011turtlebots]) through added compute resources [mattamalaNaoBackpack2017, chatzilygeroudis2014naohelmet, bergeon2019raspberry], updated sensors [chatzilygeroudis2014naohelmet, avalos2017real, da2019control, paul2021object, block2019softness, song2019communication, macias2023map], modified end-effectors [devine2016real, da2019control, chu2015robotic], and alternative software stacks [bono2024openaccessnaooan, nao_image_2025, kermorgant2025baxterros2, kumar2024kinematics, wang2016pyride, ramezani2015smooth, alas2016controlling, macias2023map].
For the NAO, researchers pursued software and hardware strategies[bono2024openaccessnaooan, nao_image_2025, mattamalaNaoBackpack2017, chatzilygeroudis2014naohelmet]. Software efforts replaced the proprietary NAOqi head-unit with Ubuntu [nao_image_2025] and ROS2 [bono2024openaccessnaooan], but these require flashing new firmware[nao_image_2025] and are limited by the NAO’s small eMMC, requiring a permanently attached USB stick for operation [nao_image_2025]. Hardware approaches preserved the factory firmware but added external compute and peripherals via short, wired connections [mattamalaNaoBackpack2017, chatzilygeroudis2014naohelmet] , enabling capabilities not supported by the stock head unit.
Similar modifications exist for other humanoid and non-humanoid robots. Prior software upgrades have replaced ROS1 with ROS2 for Baxter [kermorgant2025baxterros2], abstracted low-level ROS2 interfaces for PR2 [wang2016pyride], and introduced new inverse kinematic libraries for smoother control of both robots [kumar2024kinematics, ramezani2015smooth]. Hardware upgrades have included custom end-effectors [devine2016real, da2019control, chu2015robotic, fitter2016equipping], clothing and tactile sensors to improve human-robot hugs [block2019softness], and RGB-D cameras for teleoperation and object detection [avalos2017real, da2019control, paul2021object]. Non-humanoid robotic platforms, such as the early TurtleBots have received comparable software [song2019communication, alas2016controlling] and hardware upgrades [bergeon2019raspberry, macias2023map], extending their utility beyond their official support.
@@ III Materials and Methods
Appropriateness (AA) Flat tone Moderately engaging tone Warm, polite, and engaging tone
@@ III-A NAO Robot
We use the NAO v6 humanoid robot [aldebaran_nao_2025], a 58 cm-tall, 5.6 kg bipedal platform with 25 degrees of freedom (DOF), powered by an Intel Atom E3845 CPU with 4 GB RAM and a 32 GB eMMC storage running a Gentoo‑based Linux OS. The robot has two head-mounted HD cameras, a torso-mounted inertial measurement unit (IMU), and two ultrasonic sonars, four directional microphones, two speakers, nine capacitive touch sensors (three in the head, three per hand), and four bumper switches (two per foot). The platform supports Ethernet, Wi-Fi, and USB. Its lithium‑ion battery provides approximately 60-90 minutes of operation.
The NAO is controlled over a socket interface exposed by the proprietary NAOqi framework. Aldebaran provided a C++ library, Python bindings, and the Choregraphe GUI, though these require a local network connection to communicate with the socket interface. While the NAO v6 supports local code execution, it is restricted to Python 2.7, limiting developer flexibility [bono2024openaccessnaooan].
RobotLab further markets the NAO AI Edition (hereafter referred to as “NAO AI”) with added cloud-based conversational features [robotlab_2022]. In contrast, our work builds on the base NAO v6 with custom hardware and software upgrades, collectively referred to as the “Enhanced NAO:”
@@ III-B1 Computing and Power System
A Raspberry Pi 5 serves as the main processor, featuring a Broadcom BCM2712 quad-core 64-bit Arm Cortex-A76 (Armv8) running at 2.4 GHz [rpi5_2025]. The Raspberry Pi 5 provides USB, GPIO pins, and PCIe expansion, supporting a Hailo-8L accelerator for 13 TOPS of onboard vision-based inference [hailo20258l].
Peripheral connections include USB ports and CSI/DSI interfaces for additional sensors (e.g., Seek Thermal Camera). The system is powered by a 5 V/5 A USB-C battery pack that sustains all peripherals. Mechanical enclosures, including backpack and sensor housings, were 3D printed, with mounts adapted from open-source designs [chatzilygeroudis2014naohelmet] and custom components, as needed.
III-B2 Audio System
Audio capture is provided by the ReSpeaker with an embedded XMOS XVF-3000, offering built-in echo cancellation, noise suppression, and direction-of-arrival estimation [respeaker_2023].
III-B3 Camera System
The camera sensing suite integrates an Orbbec Gemini 2 L depth camera [orbbec_gemini] and a Seek Thermal camera (S314SPX) [seek_2024]. The Orbbec provides synchronized RGB-D streams with native OpenCV compatibility, while the Seek Thermal delivers thermography data and RGB thermal visualizations.
Figure 2: Sample camera frames produced by the enhanced camera system. A) Processed RGB and B) depth frames from the Orbbec Gemini 2 L and (C) processed thermal frame from the Seek Thermal (S314SPX) camera.
@@ III-C Software upgrades
The Raspberry Pi 5 runs Raspberry Pi OS with system nodes implemented in Node.JS, Python, and C++.
@@ III-C1 Camera Fusion
The Raspberry Pi 5 processes RGB, depth, and thermal streams locally. Frames are synchronized and passed through the Hailo-8L for human detection, face region extraction, and landmark identification in thermal and RGB data. This pipeline establishes integration of thermal and visual cues for future interaction logic. Because the Enhanced NAO’s upgraded cameras augment rather than replace the existing NAO AI camera system, they were not used in the pilot study. Their integration expands sensing capabilities (adding depth and thermal perception) (Tab. II) without interfering with the existing system. However, we include a passive vision pipeline integrating the Orbbec Gemini 2 L camera, Seek Thermal module, and the Hailo-8L for demonstration (Fig. 2).
@@ III-C2 Verbal Interaction Pipeline
Figure 3: Audio processing pipeline, from audio input via the ReSpeaker to a response audio played back through the NAO speakers.
The audio pipeline (Fig. 3) begins with the ReSpeaker, which outputs raw, processed, and echo-canceled streams. These signals are processed by the Open Embedded Audition System (ODAS) [grondin2018lightweightoptimizedsoundsource], for real-time speaker separation and sound source localization.
Deepgram [deepgram2025transcription] performs speech recognition and completion, bypassing NAO’s built-in speech-to-intent engine to enable unconstrained dialogue. Following Deepgram’s best practices, speech completion is detected when either the Endpointing feature identifies the end of natural speech in the raw audio stream or when the UtteranceEnd event indicates a sufficient silence gap in the transcript. Once a complete transcript is received, we align the word timings with ODAS outputs to produce speaker-aware transcripts.
Although Deepgram achieves a low transcription error rate, performance can degrade with low speaker volume, microphone quality, distance, or background noise. Additionally, a complete transcript does not always indicate that a participant has finished speaking. To address these challenges, we created three GPT-4o [openai2024gpt]-based LLM agents with distinct roles: a) correcting transcription errors in speaker-aware transcripts, b) detecting turn completion, and c) generating a contextually appropriate response. We used prompt engineering to structure the system and user prompts for each agent, ensuring accurate task execution and coherent, contextually relevant replies that sustained natural, open-ended dialogue.
To further reduce latency, responses are streamed immediately using Amazon Polly [AmazonPolly_2025]. Because Polly does not natively support streaming, we developed a sentence-segmentation method that sends ordered requests, enabling a “think and speak” behavior and reducing the delay between user input and robot output.
@@ III-D Pilot validation study
Figure 4: A user engaging in a multi-turn conversation with the Enhanced NAO during the pilot validation study.
Figure 5: Example of a participant conversation with the NAO AI and the Enhanced NAO.
We conducted a within-subjects, counter-balanced pilot study with eight participants (four male presenting, four female presenting) to compare interactions with the NAO AI and the Enhanced NAO. The study focused on evaluating conversational capabilities, the only functionality directly replaced and improved by the Enhanced NAO, using equivalent audio inputs and outputs. Each 15-minute session included two conditions, presented in counterbalanced order, with no compensation provided. In each condition, participants held a short conversation ($\geq 4$ turns; one turn defined as a user utterance followed by a robot response). As shown in Fig. 4, the NAO was placed on a table facing the seated participant, bringing it closer to eye level. To ensure consistency, the first turn was always a user greeting to the robot; subsequent turns were open-ended, with the option to prompt the robot for a joke, if needed. A comparison of example conversations from both conditions is shown in Fig. 5. After both conversations, each participant was verbally asked which conversation they preferred (e.g., in terms of ease and naturalness of conversation). Their responses were manually recoded by a research assistant. All sessions were video-recorded, manually transcribed, segmented into speaker turns, and timestamped for conversational quality assessment.
@@ III-E Quantifying quality of responses
We assessed conversational quality using a two-level framework adapted from LLM evaluation methods [miller2025evaluating]. The original framework included coherence, accuracy, clarity, relevance, and efficiency; we adapted it for conversational contexts by replacing accuracy with specificity, removing efficiency, and adding appropriateness and timing. This framework allows us to capture objective response quality and subjective user experience.
@@
where $W_{r}$ and $W_{u}$ denote the number of words spoken by the robot and the user, respectively. Their ratio quantifies relative talkativeness. Using Equation 3, an $R=1$ indicates perfectly balanced turn lengths, $R>1$ indicates robot-dominated speech, and $R<1$ indicates user-dominated speech. This value is averaged across the entire interaction. Because natural conversation generally involves shared participation, we considered $0.5>R>2$ an acceptable range. Values outside this range signal issues in turn-taking, such as overly strict end-of-turn detection (user-dominated conversation) or overly verbose robot responses (robot-dominated conversation), both of which can lead to a less comfortable interaction.
@@
Two evaluators (co-authors on this paper, who were not involved in the design or implementation of the Enhanced NAO) developed the coding schema before viewing any videos, to minimize expectation bias. All recordings were then coded by a primary researcher, with 20% independently coded by a secondary researcher to assess reliability. Although evaluators were aware of the condition assignments, all ratings followed standardized criteria to minimize bias.
@@ IV Results
Feature NAO AI [robotlab_2022] Enhanced NAO[aldebaran_nao_2025] with Hardware/Software Upgrades
@@
Microphones Four omnidirectional microphones; four channels at 48kHz or one channel at 16kHz. Accessible via NAOqi ALAudioDevice in 170ms chunks [aldebaran_nao_2025]. No algorithms to support ASR pipelines. Replaces existing NAO microphones
Seeed Studio ReSpeaker 4 Mic Array [respeaker_2023]
@@
Speech (TTS) Converts via NAOqi ALTextToSpeech. Supports pitch, speed, volume, pause, emphasis control, and voice effects. [aldebaran_altexttospeech_2025] Flexible with NAO’s default TTS engine, local TTS engines, or cloud-based TTS services. Tested implementation uses Amazon Polly [AmazonPolly_2025].
Speech (ASR) Via NAOqi ALSpeechRecognition. Only recognizes predefined keywords. Requires a set vocabulary and returns matches with confidence scores. Supports word spotting within longer speech. [aldebaran_asr] Flexible transcription using a dedicated ASR channel from the ReSpeaker. Tested implementation uses Deepgram to facilitate real-time speech-to-text.
@@
Static, intent-driven system focused on matching predefined intents with spoken keywords that trigger pre-determined responses. TTS engine has limited voices and lacks dynamic range. Supports open-ended dialogue with Deepgram [deepgram2025transcription], uses prompt engineering to tune an LLM (GPT-4o) [openai2024gpt] to generate relevant responses, and supports dynamic TTS voice generation via Amazon Polly [AmazonPolly_2025].